OpenStudy (anonymous):

Two men each have a mass of 90 kg. If the gravitational force between them is 8.64 x 10^-8 N, how far apart are they? G = 6.67 * 10^-11 N(m/kg)^2 I have been trying to figure it out and i can't. I have no one to show me how either

OpenStudy (matt101):

Just plug these numbers into Newton's law of universal gravitation: \[F={Gm_1m_2 \over r^2}\]

OpenStudy (anonymous):

I don't know how to do that. I am using APEX and it doesn't give a good description on how to plug them it :(

OpenStudy (matt101):

In that equation, F is the force, G is the gravitational constant, m(1) and m(2) are the masses of the two objects, and r is the distance between the objects. You have all the information given except for r, which you need to solve for. Let me know what you get!

OpenStudy (anonymous):

still nothing besides a ridiculous number. I've reread the material from APEX over and over and i still don't understand.

OpenStudy (matt101):

What number did you get?

OpenStudy (anonymous):

I don't know how to use the formula. I got 116581.709...

OpenStudy (anonymous):

I did everything wrong with the formula.

OpenStudy (matt101):

Hey! Let's back up for one second. The equation we're using is: \[F={Gm_1m_2 \over r^2}\] We want to solve for r, so it might make it easier to isolate r on one side. To do that, multiply both sides by r^2, then divide both sides by F. This will give you: \[r^2={Gm_1m_2 \over F}\] But this is solving for r^2, and we only want r! So the last thing to do is take the square root of both sides, leaving: \[r=\sqrt{Gm_1m_2 \over F}\] From the question we have the following information: F = 8.64 x 10^(-8) N G = 6.67 * 10^(-11) N(m/kg)^2 m(1) = 90 kg m(2) = 90 kg Does that help?

OpenStudy (anonymous):

Yes a lot more. Can i just plug that into my scientific calculator and solve?

OpenStudy (matt101):

You absolutely can! Do you understand the steps I followed?

OpenStudy (anonymous):

a little bit. I still can't figure it out! I got 79.07

OpenStudy (matt101):

Maybe you just made a mistake typing the numbers into your calc! You should have: \[r=\sqrt{6.67 \times 10^{-11} \times 90 \times 90 \over 8.64 \times 10^{-8}}\] Try it one more time!

OpenStudy (anonymous):

is it 2.5?

OpenStudy (matt101):

It sure is!

OpenStudy (anonymous):

Thank you!

OpenStudy (matt101):

Meaning the people are 2.5 m apart, which is a reasonable answer!